OpenStudy (anonymous):

I have been doing Partial Derivatives and have found dz/dx, dz/dy, d2z/dx2 and d2z/dy2. But I do not know how to find d2z/dxdy. Could somebody please help me?

OpenStudy (waynex):

You have \[\frac{\partial z}{\partial x}\], now take the partial derivative of that with respect to y: \[\frac{\partial\frac{\partial z}{\partial x}}{\partial y}.\] You can do this in the opposite order and get the same result. Start with y, then x.